EXERCISE 3 Consider a consumer who consumes two goods and has utility function u(x1, x2) = x2 + VX1. Income is m, the price oplease show all math work
Answer #1

is Guen: Ulility = u(2,22) = N₂ +55 Income om Þei bib - bi-Cr+t)b - _ Compensating variation is the amount we need to give to

We need to get optimal demand function is - max u Cd., ₂) set paith = m. dagrange function o L = wait H, td [m-pri-H2] differ

al. m-prev - X, 20 Substituting seqind - eq , we get 26 5xi = 0 452 from eq and (w we get one m-W | 4ks) - = 0 m . 42

thibleng uld, 12). V 17462 +/m- + m atm-14 24 m4b) - Line 45 ula, 2) = m(46) + 45. - mt1 чр 46 um, b, 1) = m + to Now, u (m,

is equal to Therefore compensation vouatio t/4p(HE). Equivalent variation is the amount that we need to take away from the co

i et - Itt-1 4p(tt) ht 4p(tt) - equwalent variation is equal to t/4p (lt) Consumer Surplus é at puce p : x = 1 at place p (it

Consumer surplus decreases by t/46 (ltat) The C, Ev and cs are equal algebraically Because the Utility function is quasilinea